OpenStudy (anonymous):

P varies directly as A and inversely as the square of N a=2,N=9,P=7

OpenStudy (anonymous):

hi harry potter =P

OpenStudy (anonymous):

I have it down to 7 =k2/81

OpenStudy (anonymous):

do I times 7 by 81?

OpenStudy (anonymous):

i think something is wrong

OpenStudy (anonymous):

possible answers are: A. 567/ 2 B. 63/ 2 C. 21/ 2 D. 14/ 81

OpenStudy (anonymous):

wouldn't my formula be p=KA/N^2

OpenStudy (anonymous):

567/2 is 283.5 not 284

OpenStudy (anonymous):

ohh i got it. it would be 567/2 !

OpenStudy (nurali):

There is one "directly" variable, A There is one "inversely" variable, N² Substituting A P = k·---- N² Substituting the complete set of data, where ALL the variables are given: P=7, N=9, A=2 2 7 = k·---- 9² Solve for k 2 7 = k·--- 81 7*81 = k·2 567 = 2k k=567/2

OpenStudy (nurali):

Sorry my bad

OpenStudy (anonymous):

lol checking your work made me get it. thank you! =)

OpenStudy (nurali):

Anytime.